40 units, 45o North of East
1
Expert's answer
2021-10-08T09:23:24-0400

Given,

Magnitude of the vector = 40 units

Angle "= 45^\circ" North of east East

Let the vector

Horizontal component "(\overrightarrow{r_x})=40 \cos (45^\circ) \hat{i}"

Vertical component "(\overrightarrow{r_y})=40\sin(45^\circ)\hat{j}"

Hence the required vector "\overrightarrow{r}=40\cos (45^\circ)\hat{i}+40\sin(45^\circ)\hat{j}"

"\Rightarrow |r|=\sqrt{(40\cos(45^\circ))^2+(40\sin (45^\circ))^2}"

"\Rightarrow |r|=\sqrt{20^2+20^2}=20\sqrt{2}"